OpenStudy (anonymous):

The sum of two numbers, a and b, is 12. Their difference is 8. Find the larger number

rvc (rvc):

a+b=12 a-b=8

OpenStudy (midhun.madhu1987):

a + b = 12 a - b = 8 add these 2 equations

OpenStudy (midhun.madhu1987):

b will be eliminated and find a.

rvc (rvc):

same ^-^ add both